OpenStudy (anonymous):

Rate of change and slope. So the question is "Find the slope of the line that passes through each pair of points" (2,2) I think it's 1/-1 am I right? or am I not suppose to write it as a fraction just like "-1"?

jimthompson5910 (jim_thompson5910):

you have (2,2) but the other point is missing

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

oops sorry its (2,2),(3,1)

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = (1-2)/(3-2) m = -1/1 m = -1 So the slope is -1 and you are 100% correct

OpenStudy (anonymous):

For questions like that, I'm always supposted to simplify and end up with one number? Not a fraction?

jimthompson5910 (jim_thompson5910):

the idea is to simplify as much as possible so you reduce as much as possible and if you end up with a fraction it's ok ex: 4/6 reduces to 2/3 10/2 reduces to 5/1 which is really 5

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

I totally understand now! Thank you so much.

jimthompson5910 (jim_thompson5910):

yw